OpenStudy (anonymous):

Evaluate lim(N->infinity) of E(k=1 to N) (2k/N^2)

OpenStudy (perl):

is this the question? $$ \Large { \lim_{n\to \infty } \sum_{k=1}^{n}~\frac{2k}{n^2} } $$

OpenStudy (anonymous):

yes, thanks

OpenStudy (anonymous):

not sure how to begin

OpenStudy (perl):

we can factor out 2/n^2 $$ \Large { \lim_{n\to \infty } \sum_{k=1}^{n}~\frac{2k}{n^2} \\\lim_{n\to \infty } \frac{2}{n^2} \cdot \sum_{k=1}^{n} k } $$

OpenStudy (perl):

now we know the sum of k , starting from k=1 to n. i can show you that separately

OpenStudy (perl):

$$ \Large { \lim_{n\to \infty } \sum_{k=1}^{n}~\frac{2k}{n^2} \\\lim_{n\to \infty } \frac{2}{n^2} \cdot \sum_{k=1}^{n} k \\\lim_{n\to \infty } \frac{2}{n^2} \cdot \frac{n(n+1)}{2} } $$

OpenStudy (perl):

now try to find the limit of that

OpenStudy (anonymous):

can you tell my why it was possible to factor out all but k initially?

OpenStudy (perl):

k stands for k=1,2,3,... up to n so we cant take that out

OpenStudy (perl):

but we can take out n^2

OpenStudy (anonymous):

so, can sigma be distributed?

OpenStudy (perl):

sorry i dont understand your question

OpenStudy (perl):

with respect to sigma, k is a variable 2/n^2 is a constant

OpenStudy (anonymous):

ok, think I understand

OpenStudy (anonymous):

cancelling the twos, we end up with (n^2 + n)/N^2 or (n + 1) / n

OpenStudy (anonymous):

and using the leading coefficient, I come up with an answer of 1, is that right?